OpenStudy (anonymous):

for f(x)=3x-7 and g(x)-2x-6 find (f o g)(4)

OpenStudy (anonymous):

(fog)= 3(2x-6)-7 then just plug in 4

OpenStudy (gorv):

(fog)(4)=f(g(4)) so first calculate g(4) which is =2*4-6=8-6=2 put value of g(4)=2 f(g(4))=f(2)=3*2-7=6-7=-1

OpenStudy (anonymous):

thanks